please answer 43 b.

(a) 2 +1-3 + 2)2 + (1 - 402 + J o 13. Find complex partial fractions for the following rational functions: 16 4+4 4
Answer #1

@ 4 74 - (2²) ² (1) 2 4 22 iz = (²+1)(7²) (E+) (+)(2-1) Applying partial fractions7th For finding C value: - -1 - Lasero cu] can ca ( <= coff) L-14wenn Izs-1BE Tai Zti Z -1 Zti . Answer is z- Zal Z zte zti i . -1z4 + 4 here 24+ 4 = (72)² + (2)² = (2²42 Z+2) (2²22 +2) | 16 16 → 2444 = (z?+27 +2) (Z2-22 +2) Applying partial fractions 164-2 Z - В [Z-(1+2)] [Z-(1-2) ] [z-(1+i) ] [2-(1-10) For O A 4 - 27 [[Z-(1+2)] [Z-(1-1)] Z=lt A = 4 - 2 z Z- (ti) z=1t i 4-2 (For B B=( 4+2 z [z-(1 + i)][z-6-1-1)] 4+2 (1-2) (na) - (-14) 49-2-2i --ti-i 2-2i 2 B = lt ti 4+2Z z²+2 7+2 z- Elti) z-(1-2) .